The density of 3.5 M aqueous solution of methanol (CH3OH) is 0.976 g / ml. What is the molarity of the solution? The molar mass of methanol is 32.4 g.


1
Expert's answer
2021-01-20T03:13:53-0500

molarity is 3.5M.

1L of the solution contains 3.5 mol of methanol and weighs 976 g

mass of methanol = 3.5*32.4 = 113.4 g

mass of water = 976 - 113.4 = 862.6 g

molality = (3.5 mol)/(0.8626 kg) = 4.0575 mol/kg